# Present address: ]Physik-Institut, Universität Zürich, Winterthurerstrasse 190, CH-8057 Zürich, Switzerland

### v2

@@ II Results and discussion
Figure 1 displays an overview of RIXS spectra for both samples, showcasing orbital and spin excitations. The common IL structure is depicted in panel (a) together with the antiferromagnetic spin lattice. Panels (b-c) report spectra taken at different transferred momenta along the [HH,0] cut of the 2D reciprocal space. The orbital ($dd$) excitations at high energy range (1-3 eV) show a small but significant dispersion as a function of the in-plane momentum, as discussed below [44, 43]. The sharp feature dispersing up to $\sim 200$ meV in PNO and $\sim 400$ meV in CCO is assigned to $\Delta S=1$ spin excitations, in agreement with previous RIXS results [12, 45, 46, 47, 48, 49, 50]. These spin excitations follow a single-magnon dispersion across most of the reciprocal space along the M-$\Gamma$-X path (panels d-e). Whereas the sharpness of the magnetic peak in CCO is a direct consequence of the long-range antiferromagnetic (AFM) order of the undoped compound [51, 50], the coexistence of superconductivity and sharp spin excitations in IL nickelates demonstrates the mild effect of self-doping on magnetism. This constitutes an important difference with respect to cuprates, where the chemical hole doping, required for superconductivity, quickly disrupts spin order and progressively smears the magnon excitation peak [49] across the superconducting dome (see below for a more quantitative comparison). Our observations are consistent with the muon spin rotation measurements in superconducting IL nickelates by Fowlie et al. [39], which hint at a similar coexistence.
@@
In order to determine the energy of the magnetic excitations, we fitted the elastic and phonon peaks with resolution-limited Gaussian functions, and the magnetic peak with the damped harmonic oscillator spectral shape, convolved with a Gaussian accounting for the experimental resolution [49] (see Figure 2(a-b)). A phonon peak at $\sim 80$ meV must be considered in the fitting of the spectra; phonon and magnon overtones are only included in the cuprate fit, because in the nickelate these features are too close in energy and too broad. A similar fit was performed on the high-energy peaks related to the orbital transitions between $3d$ levels (Panels (c-d), to be discussed in the following). More details about the fittings are available in the Supplementary Information [52]. The momentum-energy dispersions for PNO and CCO are reported in Figure 3(a), while 3(b) shows a comparison between the extracted damping coefficients, related to the spectral broadening of the magnon, and the corresponding values for superconducting (Bi,Pb)2(Sr,La)2CuO6+δ (Bi2201) samples at different doping levels, taken from Ref. 49. The peak broadening for the nickelate is found to be nearly constant ($\sim 100$ meV) over the explored momentum range: such value is obviously larger than for CCO, where long-range AFM order provides almost undamped spin wave propagation. At the same time, it is less than half the corresponding value for the lowest doping level of superconducting Bi2201 ($\gtrsim 200$ meV, within the confidence intervals). This demonstrates that self-doping has a much milder impact on spin order than chemical doping, endowing IL nickelates with a non-disruptive way to achieve superconductivity which is totally absent in copper oxides.
@@ III Methods
IV Data availability
All data shown in the main text and in the supplementary information are available at the Zenodo repository at [URL to be inserted at final submission]
Acknowledgements.
We thank Marco Moretti Sala for valuable discussion. This research used ESRF beamline ID32 under the Proposals No. HC5438, using the ERIXS spectrometer designed jointly by the ESRF and Politecnico di Milano. F.R., G.M., L.M., M.Z., M.S. and G.G. acknowledge support by the projects PRIN2017 Quantum-2D–ID 2017Z8TS5Band PRIN 2020 QT-FLUO ID 20207ZXT4Z of the Ministry for University and Research (MIUR) of Italy. D.P. acknowledges the ANR–FRANCE (French National Research Agency) for its financial support of the T-ERC project ORBIFUN ANR-23-ERCS-0003-01. This work was partially performed within the MUSA‐Multilayered Urban Sustainability Action project, funded by the European Union‐NextGenerationEU, under the National Recovery and Resilience Plan (NRRP) Mission 4 Component 2 Investment Line 1.5: Strengthening of research structures and creation of R&D,“innovation ecosystems” set up of “territorial leaders” in R&D. The work by G.M. was jointly supported by Politecnico di Milano and European X-ray Free Electron Laser Facility GmbH.
VI Author contributions
F.R., G.M., L.M., N.B., M.S. and G.G. conceived and performed the RIXS measurements, with suggestions from R.A. and D.P. H.S. and D.P. grew and characterized the PNO thin films; D.D.C. grew and characterized the CCO films. F.R. analyzed the RIXS experimental data and performed the fittings. F.R., M.S. and G.G. discussed and interpreted the results. M.Z. performed single-ion calculations. F.R. and G.G. wrote the manuscript with major suggestions from R.A. and contributions from all authors.
VII Competing interests
The authors declare no competing interests.
VIII Supplementary Information
The online version contains supplementary material available at [52]

### v3

@@ II Results and discussion
Figure 1 displays an overview of RIXS spectra for both samples, showcasing orbital and spin excitations. The common IL structure is depicted in panel (a) together with the antiferromagnetic spin lattice. Panels (b-c) report spectra taken at different transferred momenta along the [HH,0] cut of the 2D reciprocal space. The orbital ($dd$) excitations at high energy range (1-3 eV) show a small but significant dispersion as a function of the in-plane momentum, as discussed below [44, 43]. The sharp feature dispersing up to $\sim 200$ meV in PNO and $\sim 400$ meV in CCO is assigned to $\Delta S=1$ spin excitations, in agreement with previous RIXS results [12, 45, 46, 47, 48, 49, 50]. These spin excitations follow a single-magnon dispersion across most of the reciprocal space along the M-$\Gamma$-X path (panels d-e). Whereas the sharpness of the magnetic peak in CCO is a direct consequence of the long-range antiferromagnetic (AFM) order of the undoped compound [51, 50], the coexistence of superconductivity and sharp spin excitations in IL nickelates demonstrates the mild effect of self-doping on magnetism. This constitutes an important difference with respect to cuprates, where the chemical hole doping, required for superconductivity, quickly disrupts spin order and progressively smears the magnon excitation peak [49] across the superconducting dome (see below for a more quantitative comparison). Our observations are consistent with the muon spin rotation measurements in superconducting IL nickelates by Fowlie et al. [39], which hint at a similar coexistence. A recent debate was sparked by the observation of an elastic intensity increase around the quasi-commensurate wavevector $Q\sim(0.33,0)$. This was first interpreted as an evidence of charge order analogous to the Charge Density Waves observed in cuprates [52, 45, 53], but later dismissed as resulting from periodic arrangement of residual apical oxygens from incomplete topotactic reduction [54]. Our integrals of the spectral weight in the quasielastic region, displayed in Figure S6 of the Supplementary Material, show no evidence of such an increase, confirming the absence of charge order in our samples and the fully occurred topotactic reaction.
@@
In order to determine the energy of the magnetic excitations, we fitted the elastic and phonon peaks with resolution-limited Gaussian functions, and the magnetic peak with the damped harmonic oscillator spectral shape, convolved with a Gaussian accounting for the experimental resolution [49] (see Figure 2(a-b)). A phonon peak at $\sim 80$ meV must be considered in the fitting of the spectra; phonon and magnon overtones are only included in the cuprate fit, because in the nickelate these features are too close in energy and too broad. A similar fit was performed on the high-energy peaks related to the orbital transitions between $3d$ levels (Panels (c-d), to be discussed in the following). More details about the fittings are available in the Supplementary Information IV. The momentum-energy dispersions for PNO and CCO are reported in Figure 3(a), while 3(b) shows a comparison between the extracted damping coefficients, related to the spectral broadening of the magnon, and the corresponding values for superconducting (Bi,Pb)2(Sr,La)2CuO6+δ (Bi2201) samples at different doping levels, taken from Ref. 49. The peak broadening for the nickelate is found to be nearly constant ($\sim 100$ meV) over the explored momentum range: such value is obviously larger than for CCO, where long-range AFM order provides almost undamped spin wave propagation. At the same time, it is less than half the corresponding value for the lowest doping level of superconducting Bi2201 ($\gtrsim 200$ meV, within the confidence intervals). This demonstrates that self-doping has a much milder impact on spin order than chemical doping, endowing IL nickelates with a non-disruptive way to achieve superconductivity which is totally absent in copper oxides.
@@ III Methods
IV Supplementary Material
IV.1 Absence of charge order
The quasi elastic integrals of the $[H;0]$ spectra of PNO are reported in figure S6. No detectable intensity increase is present around the wavevector $\textbf{Q}=[1/3;0]$, where charge order in IL nickelates is supposed to appear[52, 45, 54].
Fig. S6: Absence of charge order. Integrals of $[H;0]$ spectra of PNO (figure 1(b) in the main text) on the quasielastic energy range [0.1;01][-0.1;0-1] eV. Error bars are given from shot noise estimation in the experimental spectra, evaluated as $\sigma\,\propto\sqrt{\text{Single Photon Counting}}$.
IV.2 RIXS spectra fitting
Low-energy RIXS fittings reported in Figure 2(a-b) in the main text were carried out by employing two Gaussian functions for the elastic and phonon peak. As explained, the magnon was fitted with a damped harmonic oscillator function [49]: we consider that here a single-peak fit is accurate enough for our scopes, although for CCO it results in an overestimation of the single-magnon damping energy close to (0.5,0). In PNO, another broad Gaussian on the right accounts for the tail of higher-energy features, mainly the lowest-energy $dd$ and the RE hybridization peak at 0.6 eV [11]. For CCO we also employed an antisymmetrized exponential starting at $0.3$ eV to take into account multiple magnetic excitations (bimagnons), and a Gaussian for the phonon overtone (biphonon). Such higher order peaks are not independently included in the PNO fit, because these features are too close in energy and too broad. Our antisymmetrization process started from a simple Gaussian:
$G(I,E,\sigma,x)=Ie^{-\frac{x-E}{2\sigma^{2}}};$ (2)
where $E,I,\sigma$ are the fitting parameters and $x$ is the energy loss variable. A symmetric copy with respect to $x=0$ was then subtracted, and the resulting odd-symmetric function was then summed to its absolute value and divided by 2. As a result, we obtain a continuous curve , that is identically null for $x\leq 0$:
$Y(I,E,\sigma,x)=\frac{G(I,E,\sigma,x)-G(I,-E,\sigma,x)+\bigg|G(I,E,\sigma,x)-G(I,-E,\sigma,x)\bigg|}{2}$ (3)
Finally, the curve was translated rightwards to make it start from 0.20.30.2-0.3 eV, which is the typical energy for multiple magnetic excitations in cuprates. The precise starting point was fixed slightly differently for different spectra, according to the quality of the resulting fit.
For fittings of the $dd$ orbital peaks, Voigt profiles antisymmetrized in a similar way were employed (Figure 2(c-d) in the main text).